# Solving Instance Detection from an Open-World Perspective

### v1

@@ 3.2.2 Foundation Model Adaptation
Hard example sampling. To construct training batches of triplets, it is important to effectively sample negative examples, overly easy negative examples may fail to provide gradients during training [51]. In this work, we adopt the batch-level hard negative sampling strategy. In each training batch, we first sample a reference image as an anchor . We then sample a random example from the same instance as the positive one . We finally sample a hard negative example from the union of the reference images of all the other instances and the distractors (explained in the next subsection). Mathematically, we find hard negative patch with the objective: , where represents the instance ID of the data , and indicates the total number of visual reference images.
3.2.3 The Proposed Data Augmentation
Novel-view synthesis. Each object instance might be represented by only a few visual examples, especially in the CID setting which provides limited real visual references for instances of interest. To augment available visual references for better adapting FM, we learn a NeRF [36] for each instance and use NeRF to synthesize novel-view images. In this process, we estimate relative camera poses (required by NeRF) using COLMAP [48], as commonly done in NeRF methods [36, 1] and InsDet approaches [27, 49]. Fig. 3 earlier displays some random NeRF-generated novel-view images with good quality in visuals.
It is worth noting that a NeRF needs to be trained only once on the visual references of an object instance. Once it is trained, we use it to generate more visual references not only for training but for testing. In particular, such NeRF-synthesized novel-view images can be stored together with the original visual references in testing for proposal-instance matching. Doing so significantly improves InsDet performance (Table 3 and Fig. 6). To the best of our knowledge, our work is the first that exploits NeRF to synthesize data for InsDet.
@@ 4.1 Experimental Setup
Compared methods. We compare the state-of-the-art methods introduced in Section 3.2.1. These methods have options to use different components (so do ours). For example, CPL can train a detector using one of the popular architectures Faster RCNN [46], RetinaNet [28], CenterNet [66], FCOS [56], and DINO [61]. OTS-FM can use SAM or GroundingDINO for proposal detection. We implement these methods by using different backbones to improve their performance (Table 1). We use subscripts to mark the backbones used in each method. Moreover, we compare more methods in the NID setting, including OS2D [40], DTOID [35], and OLN [20].
Implementation details. We carry out all experiments on a single NVIDIA 3090 GPU and implement methods with PyTorch. Unless otherwise specified, our IDOW uses GroundingDINO [30] as the proposal detector and DINOv2 [39] as the feature foundation model to finetune. When finetuning, we set hyper-parameter in the metric learning loss function. We use Adam optimizer with a learning rate 1e-3 and weight decay 0.5. We set the batch size as 100 and finetune the model for 10 epochs. We include our demo code in the supplementary material and will open-source it after the paper is accepted.
Support Ground Truth VoxDet [27] OTS-FM [49] IDOW (ours)
@@ 4.2 Benchmarking Results
Quantitative results. We evaluate our method against prior approaches on the InsDet and RoboTools dataset. Table 1 and 2 list details results in the CID and NID settings, respectively. We summarize three salient observations: First, approaching InsDet from the open-world perspective significantly outperforms previous methods. From both CID and NID settings, we find that, despite not being explicitly trained on the instances of interests, OTS-FM already achieves 10 AP than traditional detector-based methods which particularly train on such instances. Adapting FMs by our IDOW boosts performance by 3-5 AP. Second, IDOW significantly outperforms state-of-the-art InsDet methods, achieving over 5 AP higher on average than methods like OTS-FM in both CID and NID settings, demonstrating the effectiveness of our techniques. Third, IDOW is versatile and applicable to different open-world pretrained FMs. By comparing IDOW and IDOW, we find that adopting stronger FMs (GroundingDINO vs. SAM) achieves better performance.
@@ 5 Discussions
Broader Impacts. Approaching InsDet in the open-world has a significant impact on various downstream applications, such as robotics and AR/VR [64]. Detecting object instances in RGB images is considered the first step of contemporary perception algorithms. A stronger InsDet model provides richer information for the following steps, such as grasping or trajectory forecasting in robotics. Additionally, our work shows the promising performance of adapting FMs to InsDet, which will foster future research in studying different properties of FMs, such as adapting FMs to other tasks. Nonetheless, our research utilizes FMs that are potentially pretrained on private data. At this point, it is unclear how such FMs might have bias or ethical issues. As a result, we have no safeguard measure to ensure that the finetuned/adapted FMs are fair, even though we find no visible issues in the image-only dataset used for finetuning. Further, adopting SAM may result in missed detections of critical instances relevant to specific downstream applications, such as medical devices or other rare objects. Such potential issues are not addressed in this work.
@@ A Open-Source Code
We release open-source code in the form of Jupyter Notebook plus Python files.
@@ B Ablation Study and Further Analysis
Comparisons with non-Foundation Model (non-FM) methods. In Fig. 9, we use tSNE to visualize features computed by our ft-DINOv2 (in the left panel) and other non-FM methods including ft-Res101 and CPL. The middle panel shows that we use our method to finetune ImageNet-pretrained ResNet101 (ft-Res101) and the right panel shows that training a FasterRCNN with Cut-Paste-Learn [9]. We color feature points w.r.t instance IDs and overlay star-marked visual references. We can observe that ft-DINOv2 matches proposals and references better than other models. It’s worthy to note that we use the pretrained [CLS] token to extract features and class embedding to compute similarity scores for matching.
Support Ground Truth Cut-Paste-Learn [9] OTS-FM [49] IDOW (ours)
Support Ground Truth VoxDet [27] OTS-FM [49] IDOW (ours)
C Additional Visualizations
Prediction visualizations. We present more detection comparisons under the CID setting on the HR-InsDet dataset in Fig. 10 and under the NID setting on the RoboTools dataset in Fig. 11. We attach instance IDs to ground-truth and predictions to highlight whether the instance recognition is correct compared to the visual references. Our proposed method IDOW is compared with two previous arts, Cut-Paste-Learn [9] and OTS-FM [49]. We observe that our proposed method is more robust under small size, similar appearance, pose variation and serve occlusion by embracing foundation models with NeRF augmentation. For example, in Fig. 10, although instances (No.3, No.13, No.44 in the top row, and No.9, No.14, No.20, No.21, No.36, No.37 in the bottom row) in HR-InsDet are partially distracted, our IDOW can accurately identify them. In the RoboTools benchmark, most instances are very small and placed with arbitrary pose variations in the cluttered scenes. Moreover, the testing scene images are blurred. From Fig. 11, we observe that our IDOW can accurately detect more instances than VoxDet and OTS-FM although the visual references are not seen during adaptation in the NID setting.

### v2

@@ 3.2.2 The Proposed Data Augmentation
Novel-view synthesis. Each object instance might be represented by only a few visual examples, especially in the CID setting which provides limited real visual references for instances of interest. To augment available visual references for better adapting FM, we learn a NeRF [35] for each instance and use NeRF to synthesize novel-view images. In this process, we estimate relative camera poses (required by NeRF) using COLMAP [47], as commonly done in NeRF methods [35, 1] and InsDet approaches [26, 48]. Fig. 5 displays some random NeRF-generated novel-view images with good quality in visuals. It is worth noting that a NeRF needs to be trained only once on the visual references of an object instance. Once it is trained, we use it to generate more visual references not only for training but for testing. In particular, such NeRF-synthesized novel-view images can be stored together with the original visual references in testing for proposal-instance matching. Doing so significantly improves InsDet performance (Table 3 and Fig. 7). To the best of our knowledge, our work is the first that exploits NeRF to synthesize data for InsDet.
@@ 4.1 Experimental Setup
Compared methods. We compare previous methods including CPL, OTS-FM, and VoxDet. These methods have options to use different components (so do ours). For example, CPL can train a detector using one of the popular architectures Faster RCNN [45], RetinaNet [27], CenterNet [64], FCOS [54], and DINO [59]. OTS-FM can use SAM or GroundingDINO for proposal detection. We implement these methods by using different backbones to improve their performance (Table 1). We use subscripts to mark the backbones used in each method. Moreover, we compare more methods in the NID setting, including OS2D [39], DTOID [34], and OLN [20].
Implementation details. We carry out all experiments on a single NVIDIA 3090 GPU and implement methods with PyTorch. Unless otherwise specified, our IDOW uses GroundingDINO [29] as the proposal detector and DINOv2 [38] as the feature foundation model to finetune. We use the pretrained [CLS] token to extract features and class embedding to compute similarity scores for matching. When finetuning, we set hyper-parameter in the metric learning loss. We use Adam optimizer with a learning rate 1e-3 and weight decay 0.5. We set the batch size as 100 and finetune the model for 10 epochs.
@@ 4.2 Benchmarking Results
Quantitative results. We evaluate our method against prior approaches on the InsDet and RoboTools dataset. Table 1 and 2 list detailed results in the CID and NID settings, respectively. We summarize three salient observations:
First, approaching InsDet from the open-world perspective significantly outperforms previous methods. From both CID and NID settings, we find that, despite not being explicitly trained on the instances of interests, OTS-FM already achieves 10 AP than traditional detector-based methods which particularly train on such instances. Adapting FMs by our IDOW boosts performance by 3-5 AP.
Second, IDOW significantly outperforms state-of-the-art InsDet methods, achieving over 5 AP higher on average than methods like OTS-FM in both CID and NID settings, demonstrating the effectiveness of our techniques.
Third, IDOW is versatile and applicable to different open-world pretrained FMs. By comparing IDOW and IDOW, we find that adopting stronger FMs (GroundingDINO vs. SAM) achieves better performance.
Qualitative Results. Fig. 3 and 4 compares InsDet results by different methods in the CID and NID settings, respectively. We can see that our IDOW has higher precision and detects more wanted instances, which are small and under low lighting conditions. Fig. 6 uses tSNE to visualize features of visual references and detected proposals (in the HR-InsDet dataset) computed by our finetuned DINOv2 (ft-DINOv2) in the left panel, and other non-FM methods including a finetuned ImageNet-pretrained ResNet101 backbone (ft-ImageNet-Res101) and CPL in the middle and right panels, respectively. Clearly, ft-DINOv2 matches proposals and references much better than other models.
@@ 5 Discussions
Broader Impacts. Approaching InsDet in the open-world has a significant impact on various downstream applications, such as robotics and AR/VR. Detecting object instances in RGB images is considered the first step of contemporary perception algorithms. A stronger InsDet model provides richer information for the following steps, such as grasping or trajectory forecasting in robotics. Additionally, our work shows the promising performance of adapting FMs to InsDet, which will foster future research in studying different properties of FMs, such as adapting FMs to other tasks. Nonetheless, our research utilizes FMs that are potentially pretrained on private data. At this point, it is unclear how such FMs might have bias or ethical issues. As a result, we have no safeguard measure to ensure that the finetuned/adapted FMs are fair, even though we find no visible issues in the image-only dataset used for finetuning. Further, adopting SAM may result in missed detections of critical instances relevant to specific downstream applications, such as medical devices or other rare objects. Such potential issues are not addressed in this work.
@@ A Open-Source Code
In the project page (https://shenqq377.github.io/IDOW/), we release open-source code in the form of Jupyter Notebook plus Python files.
@@ B Ablation Study and Further Analysis
Support Ground Truth Cut-Paste-Learn [9] OTS-FM [48] IDOW (ours)
Support Ground Truth VoxDet [26] OTS-FM [48] IDOW (ours)
C Additional Visualizations
Prediction visualizations. We present more detection comparisons under the CID setting on the HR-InsDet dataset in Fig. 10, and under the NID setting on the RoboTools dataset in Fig. 11. We attach instance IDs to ground-truth and predictions to highlight whether the instance recognition is correct compared to the visual references. Our proposed method IDOW is compared with two previous arts, Cut-Paste-Learn [9] and OTS-FM [48]. We observe that our proposed method is more robust under small size, similar appearance, pose variation and serve occlusion by embracing foundation models with NeRF augmentation. For example, in Fig. 10, although instances (No.3, No.13, No.44 in the top row, and No.9, No.14, No.20, No.21, No.36, No.37 in the bottom row) in HR-InsDet are partially distracted, our IDOW can accurately identify them. In the RoboTools benchmark, most instances are very small and placed with arbitrary pose variations in the cluttered scenes. Moreover, the testing scene images are blurred. From Fig. 11, we observe that our IDOW can accurately detect more instances than VoxDet and OTS-FM although the visual references are not seen during adaptation in the NID setting.